OpenStudy (anonymous):

use inductive reasoning to describe each pattern 3, 8, 13, 18

OpenStudy (aum):

If I add 5 to 3 I get 8 If I add 5 to 8 I get 13 If I add 5 to 12 I get 18 ... Add 5 to the current term to get the next term.

OpenStudy (anonymous):

thanks! what about 7,4,1,-2?

OpenStudy (aum):

the last one is If I add 5 to 13 I get 18 (13 not 12)

OpenStudy (anonymous):

its an arithmaetic progression the common difference being 3.

OpenStudy (anonymous):

7-3=4 4-3=1 1-3=-2

OpenStudy (aum):

7,4,1,-2 What can be added or subtracted to go from 7 to 4? from 4 to 1?

OpenStudy (anonymous):

Subtract 3, subtract 3 @aum

OpenStudy (aum):

Yes, the pattern is: subtract 3 from the current term to get the next term.